INAF LBC+MODS queue observing, 2012 May 16-17

Observer: K. Boutsia, M. Fumana
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson
Instrument: MODS + LBC

Summary

Thick clouds on sky at the beginning of the night. We opened the dome ~4:00UT
We started with LBC but the seeing was rather bad like previous nights. At some point it started getting bettere but then we had to do the transit that requests defocusing.
At 7:20UT, while we were observing Nonino_19, dofpia started giving problems. After some consulting we realised that the disk on obs2 was full and this is why it would not work. We switched desktop and we kept observing on obs3. ~1hr lost to this problem.

TRANSIT
Because of the dofpia and full disk problems, we did not manage to start the Transit OBs at the requested time (8:13UT), but only at 8:40. During the transit the seeing was variable and the counts on the source changed from an initial 35000c to 55000c and oscilated in between. At 10:27 clouds must have started passing through, because the flux of the source became <15000c. We don't know how useful this dataset can be. Flats for this project were obtained yesterday on a blanck field as the PI requested. We did try to get some flats this morning but we are not sure if there are clouds in that direction. Try to use last night's flats first.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Radovich_B PLCK gsloan - rsloan 1hr seeing measured on images: blue:1.5" / red 1.3"
Nonino_19 EGS Bbess - rsloan 40mi seeing measured on images: blue:1.1" / red 1.1"
Nonino_19 EGS Usp - rsloan 30mi seeing measured on images: blue:1.3" / red 1.0"
Pagano GJ1214 Bbess - Rbess 3hr20min Transit - very unstable conditions - variable seeing and passing clouds during the last OB

Details

10:20 25Bias_Bino

04:07 Slew to 2012A_8B1pf
  • dofpia
  • very few stars visible... blue took some time to converge
  • lbcrangebal

04:25 2012A_8B1

04:34 Aborted because on the first image stars do not seem very well collimated
  • dofpia - again

04:38 2012A_8B1
  • dimm seeing 1.1" - airmass 1.54

05:16 Slew to Nonino *dofpia
  • dimm seeing ~1.4
  • lbcrangebal

05:32 EGS_B_R_3
  • dimm seeing ~1.1

06:00 EGS_B_R_4
  • dofpia
  • dimm seeing ~1.1

06:37 Slew to EGS_USDT_R_5
  • dofpia

07:20 dofpia (idl gave error - we need to exit and start it again)
  • this did not fix the problem - dave is looking into it

07:35 Change of configuration file for LBC

07:38 Slew to 02_OBJECT (Pagano)
  • dofpia
  • it does not work

08:00 dofpia problem was not solved but we start taking bias frames for the Pagano program
  • there is spherical on the dofpia image...

08:15 dofpia problem solved
  • we take test images to do the defocus.
  • the blue is ~50000counts and we try to defocus - something goes terribly wrong and the blue is completely out of focus
  • we run dofpia,/blueonly to fix the blue image
  • at the end of all this, we had already lost the first 25min before the transit
  • we start with the 3rd OB.

08:42 03_OBJECT

09:03 The OB was stopped

09:05 04_OBJECT
  • transit starts at 09:18
  • dimms seeing :1.2"

09:29 The OB was stopped

09:30 05_OBJECT

09:55 The OB was stopped

09:55 06_OBJECT
  • transit finishes at 10:12

10:20 The OB was stopped

10:21 07_OBJECT
  • 10:27 - clouds?? - counts went from 45000ADU to 15000ADU
  • dimm seeing:1.2"

10:41 Almost nothing vilible in the images... Clouds???

10:44 08_OBJECT - but it is rather useless... almost no stars seen in the image!!
  • stars come and go - clouds

11:12 OB aborted as requested by PI and twilight has started

11:13 09_BIAS

11:27 SkyFlatTest_BR_pa0 - just to try

11:50 End of night and RUN

-- KonstantinaBoutsia - 17 May 2012
Topic revision: r4 - 17 May 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback